OpenStudy (anonymous):

3 7/8 + 2 1/2 .. is the final answer 50/8? Thank you

OpenStudy (kainui):

Convert the numbers into complete fractions 3 7/8 = 31/8 and 2 1/2 = 5/2. Now multiply them to get the same denominator, 31/8 = 31/8 (unchanged) 5/2 = 20/8 and now you can add them together and simplify if possible, = 51/8